What is normal modes of linear triatomic molecule, when is rotating around its center of mass with angular velocity omega?
The middle molecule mass is M
and the other molecules mass is m
and the spring constant is k
normal modes of linear triatomic molecule is soleved in Herbert Goldstein's classical mechanics book, when is not rotating.
1
Expert's answer
2020-01-29T14:11:58-0500

As it is rotating with the angular velocity "\omega"

Let the radius of the molecule is r,

And due to rotation of the molecules, spring is getting elongate.

"mx\omega^2=kx"

"\Rightarrow x=\dfrac{mr\omega^2}{k}"

So, both springs will elongate "x=\dfrac{mr\omega^2}{k}"

"\omega=\sqrt{\dfrac{k}{m}}"

new angular velocity "\omega_1=\sqrt{\dfrac{k}{m}(1+\dfrac{2m}{M})}"

now, "\omega_1=\omega\sqrt{(1+\dfrac{2m}{M})}"